OpenStudy (anonymous):

the distance of a car from a certain intersection after t seconds is given by d(t)=14+4t-t^2.when(that is,for which values of t) it the car 14 units away from the refrence point ? A.t=0 B.t=2 C.t=4 D.t=6 E.never

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

can anyone help??

OpenStudy (cwrw238):

d (t) = 14 +4t - t^2 = 14 solve for t

OpenStudy (anonymous):

would it be t=4?

OpenStudy (cwrw238):

-> 4t - t^2 = 0 can u solve this for t?

OpenStudy (anonymous):

so it would be A

OpenStudy (cwrw238):

thats one value - but theres another one

OpenStudy (anonymous):

ok , so what would it be ?

OpenStudy (cwrw238):

t(4 - t)= 0 t = 0 or t = ?

OpenStudy (cwrw238):

t could be 0 but 4 - t could be zero also

OpenStudy (anonymous):

so it would be A and C

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

ok can u help with some more ?

OpenStudy (anonymous):

Which of the following are solutions to the equation below 7x^2+35x-252=0 A.4 B.-7 C.-4 D.252 E.-9 F.9